I think of a number subtract 6 multiply by 2 the anwser is 5 times the number
Answer by Cromlix(4381)

Hi there,
I think of a number subtract 6 multiply
by 2 the answer is 5 times the number.
Make number = x
(x - 6)2 = 5x
2x - 12 = 5x
Collect like terms
2x - 5x = 12
-3x = 12
x = 12/-3
x = -4
The number is -4
Hope this helps :-)